OpenStudy (anonymous):

**Please help me** [WORK SHOWN] Roberta deposited $3,116.92 into a savings account with an interest rate of 3.9% compounded quarterly. About how long will it take for the account to be worth $7,000? 14 years, 5 months 20 years, 10 months 21 years, 2 months 29 years, 11 months With the equation I did I got 21 years, 2 months. Am I correct? 3116.92(1+0.039/4)^4*21.2 = 7096.77

OpenStudy (anonymous):

7000 = 3116.92(1 + 0.039/4)^4t = 3116.92(1.00975)^4t 7000/3116.92 = 2.24581 = 1.00975^4t 4tLog 1.00975 = Log 2.24581 4t ≈ 83.385 t = 20.846 ≈ 20 years 10 months

OpenStudy (anonymous):

That's what popped up when I googled the answer

OpenStudy (anonymous):

yea why did you put that 21.2 in 3116.92(1+0.039/4)^4*21.2

OpenStudy (anonymous):

Because that's how you do the formula

OpenStudy (anonymous):

nope it isn't..you still don't get it?

OpenStudy (anonymous):

Like for example, this question Rachel deposited $6,449.82 into a savings account with an interest rate of 6.9% compounded annually. About how long will it take for the account to be worth $8,000? 3 years, 3 months 10 years, 10 months 12 years, 0 months 2 years, 4 months

OpenStudy (anonymous):

6449.821+0.039/1)^1*3.3 = 8038.48

OpenStudy (anonymous):

I used the same formula, the answer I chose was 3 years, 3 months and it was correct.

OpenStudy (anonymous):

from where did that 3.3 pop up?

OpenStudy (anonymous):

3 years, 3 months

OpenStudy (anonymous):

means you are using guessed answer to calculate answer?

OpenStudy (anonymous):

No? I'm using the formula given by my teacher to find the answer.

OpenStudy (anonymous):

what if you didn't know 3.3 was the answer?

OpenStudy (anonymous):

I did because I calculated EVERY single answer choice there, 3 years, 3 months WHICH COVERTED TO 3.3 was the only one over $8,000

OpenStudy (anonymous):

you have to put t instead of 3.3 and then solve for t

OpenStudy (anonymous):

10 years, 10 months = 10.10 12 years, 0 months = 12 2 years, 4 months =2.4

OpenStudy (sumi29):

@stgreen has already told you the answer, but in general for most questions involving multiple choices, its better to find an equation and solve for the variable rather than plugging in the options as its far less time-consuming.

OpenStudy (anonymous):

you calculated every single ans?that's laborious..you could've put t there and got the correct one in a single try

OpenStudy (anonymous):

Well this was the formula given to me

OpenStudy (anonymous):

@sumi29 explained my point

OpenStudy (anonymous):

And I got the answer right using this formula

OpenStudy (anonymous):

Oh well, whatever

OpenStudy (sumi29):

@murrcat You can always get the right answer plugging in the options, but imagine for a moment, what happens if you pick two or three random choices out of four and all of them turn out to be wrong? You would have simply wasted your time, time that could've been well spent on constructing an equation and solving it for your answer. Your method is not wrong, its simply not worth the effort. I hope you get it.

OpenStudy (anonymous):

Mhm